Wheat is produced under perfectly competitive conditions. Individual wheat farmers have U-shaped, long-run average cost curves that reach a minimum average cost of $3 per bushel when 1,000 bushels are products

a) If the market demand curve for wheat is given by QD = 2,600,000 - 200,000P, where Qp is the number of bushels demanded per year and P is the price per bushel, in long-run equilibrium what will be the price of wheat, how much total wheat will be demanded, and how many wheat farms will there be?
1
Expert's answer
2019-07-18T08:13:44-0400

a) If the market demand curve for wheat is given by QD = 2,600,000 - 200,000P, in long-run equilibrium the price of wheat will be P = minLATC = $3, total wheat demanded will be q = 1000 bushels, and the number of wheat farms there will be:

Qd/Q = (2,600,000 - 200,000×3)/1,000 = 2,000.